OpenStudy (anonymous):

A sled carrying 3 people is pulled by 2 ropes. One rope is pulled 215N [E13N] and the other 194N [S40E]. Determine acceleration of the sled if it has a mass of 21kg and each person weighs 980N.

OpenStudy (anonymous):

a = F/m 980N * 3 = 2940 2940/21 = 140m/s^2 Is this right?

OpenStudy (anonymous):

no. First draw a free body diagram. Using the rules for vector addition, find the magnitude and direction of the unbalanced force. Then use Newton's second law. You have the weights of the passengers. Remember to use their masses.

OpenStudy (anonymous):

Remember that the net force is given as follows:\[F _{net}=\sum_{1}^{k}F _{i}\]. In other words the net force on the sled is the sum of the individual forces acting on the sled.